# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v2

@@ 4.2 Analysis of the feasibility correction
Throughout this subsection and the remainder of this section, we assume that Assumptions 1, 2, and 3 hold, and $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$. The sequences $\{(x^{k},y^{k},\tilde{y}^{k})\}$ and $\{p_{k}\}$ are generated by AGILS (Algorithm 1), where the step sizes satisfy $\alpha_{k}\in[\underline{\alpha},2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[\underline{\beta},2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $\underline{\alpha},\underline{\beta},c_{\alpha},c_{\beta}$ are positive constants. The inexactness parameters satisfy $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$ and $\tau_{k}:=\tau_{0}/(k+1)^{p_{\tau}}$ with $\tau_{0}>0$ and $p_{\tau}>0$.
@@
Summing (30) from Lemma 5 over $k=0,1,\dots,K-1$ yields,
@@
Since $\{\zeta_{k}\}$ is square summable from Lemma 5, it follows that
@@
From Lemma 6, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\|=0$. As $p_{k_{j}}\geq p_{0}$, the above inequality ensures that $\lim_{k_{j}\to\infty}\|\tilde{r}^{k_{j}}\|=0$.
@@
Let $(\bar{x},\bar{y})$ be an accumulation point of $\{(x^{k},{y}^{k})\}$, and let $\{k_{j}\}_{j\in\mathbb{N}}\subseteq\mathbb{N}$ be an index subsequence such that $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ as $j\rightarrow\infty$. By Lemma 6, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}+1},y^{k_{j}+1})-(x^{k_{j}},\tilde{y}^{k_{j}})\|=0$, which implies that $(x^{k_{j}},\tilde{y}^{k_{j}})$ also converges to $(\bar{x},\bar{y})$
@@
The condition $\tilde{y}^{k_{j}}=y^{k_{j}}$ combined with $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$ implies that $(x^{k_{j}},{y}^{k_{j}})\to(\bar{x},\bar{y})$. From Lemma 6, we know $\lim_{k_{j}\to\infty}\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\|=0$. For the condition of Case I to hold, we must therefore have $t^{k_{j}}\rightarrow 0$. This implies
@@
Since $\zeta_{k}\rightarrow 0$ and $\mathcal{G}(\theta^{k_{j}},x^{k_{j}},y^{k_{j}})\leq\zeta_{k_{j}}$, Lemma 2 ensures that $\|\theta^{k_{j}}-\theta^{*}_{\gamma}(x^{k_{j}},y^{k_{j}})\|\rightarrow 0$. By the continuity of $\theta^{*}_{\gamma}(x,y)$ (Proposition 2), it follows that $\theta^{k_{j}}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. The continuity of $\varphi$ then implies that $\varphi(x^{k_{j}},\theta^{k_{j}})+\frac{1}{2\gamma}\|\theta^{k_{j}}-y^{k_{j}}\|^{2}\rightarrow v_{\gamma}(\bar{x},\bar{y})$. Then taking the limit in the inequality above as $j\rightarrow\infty$ yields $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon\leq 0$.
@@
In this scenario, the feasibility correction procedure is applied at each iteration $k=k_{j}-1$ to generate $\tilde{y}^{k_{j}}$ satisfying condition (24). Since $(x^{k_{j}},\tilde{y}^{k_{j}})\rightarrow(\bar{x},\bar{y})$ and satisfies condition (24), it follows from Lemma 7 that $\bar{y}\in S(\bar{x})$. For any such point, $\theta^{*}_{\gamma}(\bar{x},\bar{y})=\bar{y}$ (see, e.g., [7, Theorem 10.7]), and therefore $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})=0$. Since $\epsilon>0$, the constraint $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon$ is satisfied.
@@
Since $\{(x^{k},y^{k})\}$ is bounded and Lemma 6 establishes that $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\|=0$, the sequence $\{\tilde{y}^{k}\}$ is also bounded. By passing to a subsequence if necessary, we can assume that for the indices $\{k_{j}\}$, we have $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ for some $(\bar{x},\bar{y})\in X\times Y$ and $\tilde{y}^{k_{j}+1}\to\tilde{y}$ for some $\tilde{y}\in Y$. From Lemma 6, it also follows that $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$
@@
We will show that this leads to the conclusion that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon$. Let us assume the contrary, i.e., $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon<0$. From Lemma 2, the conditions $\mathcal{G}(\theta^{k_{j}+1},x^{k_{j}+1},y^{k_{j}+1})\leq\zeta_{k_{j}+1}$ and $\zeta_{k}\rightarrow 0$ ensure that $\|\theta^{k_{j}+1}-\theta^{*}_{\gamma}(x^{k_{j}+1},y^{k_{j}+1})\|\rightarrow 0$. By the continuity of $\theta^{*}_{\gamma}(x,y)$ (Proposition 2) and $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$, we have $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. The continuity of $\varphi$ then implies that $\varphi(x^{k_{j}+1},\theta^{k_{j}+1})+\|\theta^{k_{j}+1}-y^{k_{j}+1}\|^{2}/(2\gamma)\rightarrow v_{\gamma}(\bar{x},\bar{y})$. Our assumption $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon<0$, combined with the continuity of $\varphi$, means that for $j$ sufficiently large,
@@
Case I: $\left\|y^{k+1}-\theta^{k+1}\right\|<c_{y}\gamma/p_{k}$.
Case II: $\left\|y^{k+1}-\theta^{k+1}\right\|\geq c_{y}\gamma/p_{k}$, the feasibility correction procedure is applied, but the corrected iterate $\tilde{y}^{k+1}$ violates the descent condition (25).
@@
$\left\|y^{k_{j}+1}-\theta^{k_{j}+1}\right\|<\frac{\gamma}{p_{k_{j}+1}}\to 0.$
Since we assume $p_{k_{j}}\rightarrow\infty$, the right-hand side converges to 00. As $j\to\infty$, we have $y^{k_{j}+1}\rightarrow\bar{y}$ and $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Taking the limit yields $\|\bar{y}-\theta^{*}_{\gamma}(\bar{x},\bar{y})\|\leq 0$, which implies $\bar{y}=\theta^{*}_{\gamma}(\bar{x},\bar{y})$. By definition, this yields $\varphi(\bar{x},\bar{y})=v_{\gamma}(\bar{x},\bar{y})$, which contradicts our established result that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon>0$.
@@
As $j\rightarrow\infty$, we have $(x^{k_{j}+1},\tilde{y}^{k_{j}+1})\rightarrow(\bar{x},\tilde{y})$. Since $(x^{k_{j}+1},\tilde{y}^{k_{j}+1})$ satisfies condition (24), it follows from Lemma 7 that $\tilde{y}\in S(\bar{x})$. Moreover, given that $\zeta_{k}\rightarrow 0$ and $\mathcal{G}(\tilde{\theta}^{k_{j}+1},x^{k_{j}+1},\tilde{y}^{k_{j}+1})\leq\zeta_{k_{j}+1}$, Lemma 2 ensures that $\|\tilde{\theta}^{k_{j}+1}-\theta^{*}_{\gamma}(x^{k_{j}+1},\tilde{y}^{k_{j}+1})\|\rightarrow 0$, and thus $\tilde{\theta}^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\tilde{y})$ from the continuity of $\theta^{*}_{\gamma}(x,y)$ (Proposition 2).
@@
Because $\tilde{y}\in S(\bar{x})$, we know that $\theta^{*}_{\gamma}(\bar{x},\tilde{y})=\tilde{y}$ (see, e.g., [7, Theorem 10.7]). Taking the limit of the left-hand side of the inequality above as $k_{j}\rightarrow\infty$, and using the continuity of $F$ and $\varphi$ along with $p_{k_{j}}\rightarrow\infty$, we find that the expression converges to $\varphi(\bar{x},\tilde{y})-\varphi(\bar{x},\theta^{*}_{\gamma}(\bar{x},\tilde{y}))-\frac{1}{2\gamma}\|\theta^{*}_{\gamma}(\bar{x},\tilde{y})-\tilde{y}\|^{2}=0$. On the other hand, we have $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ and $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Therefore, the right-hand side of the inequality converges to $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon>0$. As $k_{j}\rightarrow\infty$, the inequality thus implies $00\geq\epsilon>$ 0, which is a contradiction.
@@
Since both possible scenarios for an infinite number of penalty updates lead to a contradiction, our initial assumption that $p_{k}\to\infty$ must be false. Therefore, the sequence of penalty parameters $\{p_{k}\}$ is bounded.

### v3

@@ 4.2 Analysis of the feasibility correction
Throughout this subsection and the remainder of this section, we assume that Assumptions 2.1, 2.2, and 2.3 hold, and $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$. The sequences $\{(x^{k},y^{k},\tilde{y}^{k})\}$ and $\{p_{k}\}$ are generated by AGILS (Algorithm 1), where the step sizes satisfy $\alpha_{k}\in[\underline{\alpha},2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[\underline{\beta},2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $\underline{\alpha},\underline{\beta},c_{\alpha},c_{\beta}$ are positive constants. The inexactness parameters satisfy $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$ and $\tau_{k}:=\tau_{0}/(k+1)^{p_{\tau}}$ with $\tau_{0}>0$ and $p_{\tau}>0$.
@@
Summing (30) from Lemma 4.4 over $k=0,1,\dots,K-1$ yields,
@@
Since $\{\zeta_{k}\}$ is square summable from Lemma 4.4, it follows that
@@
From Lemma 4.5, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\|=0$. As $p_{k_{j}}\geq p_{0}$, the above inequality ensures that $\lim_{k_{j}\to\infty}\|\tilde{r}^{k_{j}}\|=0$.
@@
Let $(\bar{x},\bar{y})$ be an accumulation point of $\{(x^{k},{y}^{k})\}$, and let $\{k_{j}\}_{j\in\mathbb{N}}\subseteq\mathbb{N}$ be an index subsequence such that $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ as $j\rightarrow\infty$. By Lemma 4.5, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}+1},y^{k_{j}+1})-(x^{k_{j}},\tilde{y}^{k_{j}})\|=0$, which implies that $(x^{k_{j}},\tilde{y}^{k_{j}})$ also converges to $(\bar{x},\bar{y})$ .
@@
The condition $\tilde{y}^{k_{j}}=y^{k_{j}}$ combined with $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$ implies that $(x^{k_{j}},{y}^{k_{j}})$
$\to(\bar{x},\bar{y})$. From Lemma 4.5, we know $\lim_{k_{j}\to\infty}\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\|=0$. For the condition of Case I to hold, we must therefore have $t^{k_{j}}\rightarrow 0$. This implies
@@
Since $\zeta_{k}\rightarrow 0$ and $\mathcal{G}(\theta^{k_{j}},x^{k_{j}},y^{k_{j}})\leq\zeta_{k_{j}}$, Lemma 4.1 ensures that $\|\theta^{k_{j}}-\theta^{*}_{\gamma}(x^{k_{j}},y^{k_{j}})\|\rightarrow 0$. By the continuity of $\theta^{*}_{\gamma}(x,y)$ (Proposition 2.8), it follows that $\theta^{k_{j}}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. The continuity of $\varphi$ then implies that $\varphi(x^{k_{j}},\theta^{k_{j}})+\frac{1}{2\gamma}\|\theta^{k_{j}}-y^{k_{j}}\|^{2}\rightarrow v_{\gamma}(\bar{x},\bar{y})$. Then taking the limit in the inequality above as $j\rightarrow\infty$ yields $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon\leq 0$.
@@
In this scenario, the feasibility correction procedure is applied at each iteration $k=k_{j}-1$ to generate $\tilde{y}^{k_{j}}$ satisfying condition (24). Since $(x^{k_{j}},\tilde{y}^{k_{j}})\rightarrow(\bar{x},\bar{y})$ and satisfies condition (24), it follows from Lemma 4.6 that $\bar{y}\in S(\bar{x})$. For any such point, $\theta^{*}_{\gamma}(\bar{x},\bar{y})=\bar{y}$ (see, e.g., [7, Theorem 10.7]), and therefore $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})=0$. Since $\epsilon>0$, the constraint $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon$ is satisfied.
@@
Since $\{(x^{k},y^{k})\}$ is bounded and Lemma 4.5 establishes that $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})$
$(x^{k},\tilde{y}^{k})\|=0$, the sequence $\{\tilde{y}^{k}\}$ is also bounded. By passing to a subsequence if necessary, we can assume that for the indices $\{k_{j}\}$, we have $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ for some $(\bar{x},\bar{y})\in X\times Y$ and $\tilde{y}^{k_{j}+1}\to\tilde{y}$ for some $\tilde{y}\in Y$. From Lemma 4.5, it also follows that $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$ .
@@
We will show that this leads to the conclusion that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon$. Let us assume the contrary, i.e., $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon<0$. From Lemma 4.1, the conditions $\mathcal{G}(\theta^{k_{j}+1},x^{k_{j}+1},y^{k_{j}+1})\leq\zeta_{k_{j}+1}$ and $\zeta_{k}\rightarrow 0$ ensure that $\|\theta^{k_{j}+1}-\theta^{*}_{\gamma}(x^{k_{j}+1},y^{k_{j}+1})\|\rightarrow 0$. By the continuity of $\theta^{*}_{\gamma}(x,y)$ (Proposition 2.8) and $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$, we have $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. The continuity of $\varphi$ then implies that $\varphi(x^{k_{j}+1},\theta^{k_{j}+1})+\|\theta^{k_{j}+1}-y^{k_{j}+1}\|^{2}/(2\gamma)\rightarrow v_{\gamma}(\bar{x},\bar{y})$. Our assumption $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon<0$, combined with the continuity of $\varphi$, means that for $j$ sufficiently large,
@@
Case I: $\left\|y^{k+1}-\theta^{k+1}\right\|\leq c_{y}\gamma/p_{k}$.
Case II: $\left\|y^{k+1}-\theta^{k+1}\right\|>c_{y}\gamma/p_{k}$, the feasibility correction procedure is applied, but the corrected iterate $\tilde{y}^{k+1}$ violates the descent condition (25).
@@
$\left\|y^{k_{j}+1}-\theta^{k_{j}+1}\right\|\leq\frac{\gamma}{p_{k_{j}+1}}\to 0.$
Since we assume $p_{k_{j}}\rightarrow\infty$, the right-hand side converges to 0. As $j\to\infty$, we have $y^{k_{j}+1}\rightarrow\bar{y}$ and $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Taking the limit yields $\|\bar{y}-\theta^{*}_{\gamma}(\bar{x},\bar{y})\|\leq 0$, which implies $\bar{y}=\theta^{*}_{\gamma}(\bar{x},\bar{y})$. By definition, this yields $\varphi(\bar{x},\bar{y})=v_{\gamma}(\bar{x},\bar{y})$, which contradicts our established result that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon>0$.
@@
[…]
@@
Because $\tilde{y}\in S(\bar{x})$, we know that $\theta^{*}_{\gamma}(\bar{x},\tilde{y})=\tilde{y}$ (see, e.g., [7, Theorem 10.7]). Taking the limit of the left-hand side of the inequality above as $k_{j}\rightarrow\infty$, and using the continuity of $F$ and $\varphi$ along with $p_{k_{j}}\rightarrow\infty$, we find that the expression converges to $\varphi(\bar{x},\tilde{y})-\varphi(\bar{x},\theta^{*}_{\gamma}(\bar{x},\tilde{y}))-\frac{1}{2\gamma}\|\theta^{*}_{\gamma}(\bar{x},\tilde{y})-\tilde{y}\|^{2}=0$. On the other hand, we have $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ and $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Therefore, the right-hand side of the inequality converges to $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon>0$. As $k_{j}\rightarrow\infty$, the inequality thus implies $0\geq\epsilon>0$, which is a contradiction.
@@
Since both possible scenarios for an infinite number of penalty updates lead to a contradiction, our initial assumption that $p_{k}\to\infty$ must be false. Therefore, the sequence of penalty parameters $\{p_{k}\}$ is bounded. ∎
Next, we establish that once the penalty parameter $p_{k}$ becomes sufficiently large, the feasibility correction procedure will be applied only finitely many times.
Proposition 4.9.
Let $\epsilon>0$. Suppose the sequence of iterates $\left\{(x^{k},y^{k})\right\}$ is bounded, and define $M_{F}:=\sup_{k}|F(x^{k},{y}^{k})|$. If there exists $k_{0}\in\mathbb{N}$ such that the penalty parameter $p_{k_{0}}$ is sufficiently large to satisfy $p_{k_{0}}\geq 4M_{F}/\epsilon$, then the feasibility correction procedure is applied only a finite number of times.
Proof.
By Theorem 4.8, the penalty parameter sequence $\{p_{k}\}$ is bounded. Hence, $p_{k}=\bar{p}$ for all sufficiently large $k$, , and we may assume without loss of generality that $p_{k}=\bar{p}$ for all $k$. Under the assumption, we have
$\frac{M_{F}}{\bar{p}}\leq\frac{\epsilon}{4}.$ (46)
From (29) in Lemma 4.3, together with (35) and (40) in the proof of Lemma 4.4, we obtain
$\tilde{\psi}_{\bar{p}}\left(x^{k+1},\tilde{y}^{k+1}\right)\leq$ $\tilde{\psi}_{\bar{p}}(x^{k+1},y^{k+1})+C_{\gamma}\zeta_{k+1}^{2}$ (47)
$\leq$ $\tilde{\psi}_{\bar{p}}\left(x^{k},\tilde{y}^{k}\right)+C_{1}\zeta_{k}^{2}+C_{2}\zeta_{k+1}^{2}-\frac{c_{\alpha}}{4}\|x^{k+1}-x^{k}\|^{2}-\frac{c_{\beta}}{4}\|y^{k+1}-\tilde{y}^{k}\|^{2},$
where $C_{1}=C_{\eta}^{2}(L_{f_{x}}+L_{g_{2}})^{2}/c_{\alpha}$ and $C_{2}=C_{\eta}^{2}/(c_{\beta}\gamma^{2})+C_{\gamma}$.
Define
$A_{k}:=\tilde{\psi}_{\bar{p}}(x^{k},\tilde{y}^{k})+\sum_{j=k}^{\infty}\!\big(C_{1}\zeta_{j}^{2}+C_{2}\zeta_{j+1}^{2}\big).$
Adding the tail sum to both sides of (47) yields
$A_{k+1}\leq A_{k}-\frac{c_{\alpha}}{4}\|x^{k+1}-x^{k}\|^{2}-\frac{c_{\beta}}{4}\|y^{k+1}-\tilde{y}^{k}\|^{2}.$
Thus, $\{A_{k}\}$ is a nonincreasing sequence. Since $\tilde{\psi}_{\bar{p}}\left(x^{k},\tilde{y}^{k}\right)\geq 0$ for all $k$, we have $A_{k}\geq 0$, implying that $A_{k}$} is bounded below and hence convergent, i.e., $A_{k}\to\bar{A}\in\mathbb{R}$.
From Lemma 4.4, the sequence $\{\zeta_{k}\}$ is square summable, so $\zeta_{k}\to 0$ and $\sum_{j=k}^{\infty}(C_{1}\zeta_{j}^{2}+C_{2}\zeta_{j+1}^{2})\to 0$ as $k\to\infty$. Consequently,
$\lim_{k\rightarrow\infty}\tilde{\psi}_{\bar{p}}(x^{k},\tilde{y}^{k})=\lim_{k\rightarrow\infty}\left(A_{k}-\sum_{j=k}^{\infty}\big(C_{1}\zeta_{j}^{2}+C_{2}\zeta_{j+1}^{2}\big)\right)=\bar{A}.$
Combining this with (47) and the fact that $\zeta_{k}\to 0$ gives
$\lim_{k\rightarrow\infty}\tilde{\psi}_{\bar{p}}(x^{k},y^{k})=\lim_{k\rightarrow\infty}\tilde{\psi}_{\bar{p}}(x^{k},\tilde{y}^{k})=\bar{A}.$
We now prove the claim by contradiction. Suppose that the feasibility correction procedure is triggered infinitely many times. Then there exists an infinite index set $k_{j}$} such that Case 3 is executed at each iteration $k_{j}$. Then, we have, for all $j>0$,
$t^{k_{j}+1}>0.$
By the definition of $t^{k}$, we have
$t^{k_{j}+1}$ $=\varphi(x^{k_{j}+1},y^{k_{j}+1})-\varphi(x^{k_{j}+1},\theta^{k_{j}+1})-\frac{1}{2\gamma}\|\theta^{k_{j}+1}-y^{k_{j}+1}\|^{2}-\epsilon$
$\leq\varphi(x^{k_{j}+1},y^{k_{j}+1})-v_{\gamma}(x^{k_{j}+1},y^{k_{j}+1})-\epsilon,$
where the inequality follows from $v_{\gamma}(x^{k_{j}+1},y^{k_{j}+1})=\min_{\theta\in Y}\{\varphi(x^{k_{j}+1},\theta)+\tfrac{1}{2\gamma}\|\theta-y^{k_{j}+1}\|^{2}\}$. Since $t^{k_{j}+1}>0$, we obtain
$\varphi(x^{k_{j}+1},y^{k_{j}+1})-v_{\gamma}(x^{k_{j}+1},y^{k_{j}+1})\geq\epsilon.$ (48)
As the feasibility correction procedure is performed at iteration $k_{j}$, a corrected iterate $\tilde{y}^{k_{j}+1}$ satisfying condition (24) is generated. Because $\{(x^{k},y^{k})\}$ is bounded and Lemma 4.5 shows that $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\|=0$, the sequence $\{\tilde{y}^{k}\}$ is also bounded. By passing to a subsequence if necessary, we may assume that $(x^{k_{j}+1},\tilde{y}^{k_{j}+1})\to(\bar{x},\bar{y})$ for some $(\bar{x},\bar{y})\in X\times Y$. Then Lemma 4.6 implies that $\bar{y}\in S(\bar{x})$, and thus $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})=0$. By continuity of $\varphi$ and $v_{\gamma}$, there exists $j_{0}\in\mathbb{N}$ such that for all $j\geq j_{0}$,
$\varphi(x^{k_{j}+1},\tilde{y}^{k_{j}+1})-v_{\gamma}(x^{k_{j}+1},\tilde{y}^{k_{j}+1})\leq\frac{\epsilon}{10}.$ (49)
Furthermore, since $\{(x^{k},y^{k})\}$ is bounded and $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\|=0$ from Lemma 4.5, by the continuity of $F$ on $X\times Y$, we have
$\sup_{k}|F(x^{k},\tilde{y}^{k})|=\sup_{k}|F(x^{k},{y}^{k})|=M_{F}.$ (50)
We now consider the difference $\tilde{\psi}_{\bar{p}}(x^{k_{j}+1},y^{k_{j}+1})-\tilde{\psi}_{\bar{p}}(x^{k_{j}+1},\tilde{y}^{k_{j}+1})$ when $j\geq j_{0}$,
$\tilde{\psi}_{\bar{p}}(x^{k_{j}+1},y^{k_{j}+1})-\tilde{\psi}_{\bar{p}}(x^{k_{j}+1},\tilde{y}^{k_{j}+1})$
$=$ $\frac{1}{\bar{p}}\big(F(x^{k_{j}+1},y^{k_{j}+1})-F(x^{k_{j}+1},\tilde{y}^{k_{j}+1})\big)$
$+(\varphi-v_{\gamma})(x^{k_{j}+1},y^{k_{j}+1})-(\varphi-v_{\gamma})(x^{k_{j}+1},\tilde{y}^{k_{j}+1})$
$\geq$ $-\frac{2M_{F}}{\bar{p}}+\epsilon-\frac{\epsilon}{10}$
$\geq$ $-\frac{\epsilon}{2}+\epsilon-\frac{\epsilon}{10}$
$=$ $\frac{2}{5}\epsilon>0,$
where the first inequality follows from (48), (49) and (50) and the second follows from (46). This contradicts the fact that $\lim_{k\rightarrow\infty}\tilde{\psi}_{\bar{p}}(x^{k},y^{k})=\lim_{k\rightarrow\infty}\tilde{\psi}_{\bar{p}}(x^{k},\tilde{y}^{k})=\bar{A}$. Hence, the feasibility correction procedure can occur only finitely many times. ∎